A baby weighs 9 pounds at birth and weighs 15 pounds six months later. Find the percentage increase in weight
Found 2 solutions by checkley77, Mathtut:
Answer by checkley77(12844)

15/9=1.66 or 166.66% increase.

Answer by Mathtut(3670)

The increase is 15-9=6
:
so you divide the increase by the original weight to get the percentage of increase
:
6/9=.6667(100)=66.67% increase
:
It is quite common to make the mistake of dividing the entire new weight by the original weight, but we are looking for the percentage INCREASE here.